# Forgotten Polygons: Multimodal Large Language Models are Shape-Blind

### v1

@@ Abstract
William Rudman⋆1, Michal Golovanesky⋆1, Amir Bar2, Vedant Palit3, Yann LeCun4, Carsten Eickhoff5, Ritambhara Singh1 1Brown University, 2Tel Aviv University, 3IIT Kharagpur, 4New York University, 5University of Tübingen {william_rudman, michal_golovanesky}@brown.edu
@@ 3.4 Visually-Cued CoT
| Image Labels | Molmo | Janus-Pro | GPT-4-Turbo | GPT-4o |
| Heptagon: Direct / VC-CoT | ||||
| Numbers 1-7 | 60% / 69% | 1% / 100% | 79% / 98% | 43% / 100% |
| Random Nums | 17% / 29% | 3% / 90% | 1% / 94% | 31% / 100% |
| Letters A-G | 63% / 75% | 15% / 97% | 99% / 100% | 59% / 100% |
| Random Letters | 49% / 89% | 7% / 74% | 7% / 82% | 7% / 93% |
| No Annotations | 1% / 1%† | 4% / 14%† | 1%/ 3%† | 22% / 41%† |
| Arrow on Plus: Direct / VC-CoT | ||||
| Numbers 1-15 | 0% / 61% | 13% / 92% | 97% / 99% | 88% / 99% |
| Random Nums | 0% / 5% | 0% / 0% | 1% / 78% | 3% / 77% |
| Letters A-O | 0% / 1% | 0% / 2% | 75% / 74% | 14% / 99% |
| Random Letters | 0% / 2% | 0% / 1% | 0% / 25% | 3% / 96% |
| No Annotations | 0% / 0%† | 0% / 0%† | 0% / 1%† | 0% / 1%† |
@@
Table 4 shows the performance of Molmo, Janus-Pro, GPT-4-Turbo, and GPT-4o, and Appendix Section G shows results for all other models. Molmo and Janus-Pro were explicitly designed for CoT reasoning through dedicated training, while GPT models are well-documented to excel with CoT prompting Wei et al. (2023), making them the most relevant candidates for VC-CoT. In Table 4, we see that these models exhibit substantial improvements when visually cued CoT prompts are paired with explicit visual annotations, such as numbering or labeling each side. This is especially evident in the GPT models, where GPT-4o increases from 7% to 93%, in the random letters annotation.
@@ Appendix G Visually-Cued Chain-of-Thought Shape Results
| Numbers 1-7 | 0% | 1% | 0% | 5% | 0% | 17% | 0% | 4% | 26% | 68% | 100% | 100% | 0% | 0% | 79% | 81% | 0% | 0% |
| Random Numbers | 0% | 0% | 0% | 0% | 0% | 0% | 0% | 4% | 1% | 21% | 32% | 50% | 0% | 0% | 2% | 0% | 0% | 1% |
| Letters A-G | 0% | 0% | 0% | 0% | 0% | 0% | 0% | 3% | 1% | 2% | 45% | 96% | 0% | 0% | 0% | 0% | 0% | 0% |
| Random Letters | 0% | 0% | 0% | 0% | 0% | 0% | 0% | 3% | 0% | 3% | 3% | 58% | 0% | 0% | 0% | 0% | 0% | 1% |
| No Annotations | 0% | 0% | 0% | 0% | 0% | 0% | 0% | 3% | 0% | 0% | 0% | 0% | 0% | 0% | 0% | 0% | 0% | 0% |
| Numbers 1-15 | 0% | 8% | 12% | 28% | 26% | 26% | 0% | 0% | 0% | 14% | 99% | 99% | 0% | 0% | 0% | 0% | 0% | 61% |
| Random Numbers | 0% | 0% | 0% | 0% | 0% | 0% | 0% | 0% | 0% | 3% | 32% | 21% | 0% | 0% | 0% | 0% | 0% | 28% |
| Random Letters | 0% | 0% | 0% | 0% | 0% | 0% | 0% | 1% | 0% | 1% | 0% | 84% | 0% | 0% | 0% | 0% | 0% | 0% |
@@ Appendix H VC-CoT Results on MathVerse
Despite these challenges, our results show that VC-CoT improves overall performance across these models, with an average accuracy increase from 34.51% (Direct) to 37.32% (VC-CoT). Notably, 9 out of 13 models (including ones in Table 5) benefit from visually guided reasoning, suggesting that explicit extraction and processing of visual information can help compensate for the weaknesses of generic CoT prompting.

### v2

@@ Abstract
William Rudman⋆1, Michal Golovanevsky⋆1, Amir Bar2, Vedant Palit3, Yann LeCun4, Carsten Eickhoff5, Ritambhara Singh1 1Brown University, 2Tel Aviv University, 3IIT Kharagpur, 4New York University, 5University of Tübingen {william_rudman, michal_golovanesky}@brown.edu
@@ 3.4 Visually-Cued CoT
| Image Labels | Qwen2-VL | Molmo | Janus-Pro | GPT-4-Turbo | GPT-4o |
| Heptagon: Direct / VC-CoT | |||||
| Numbers 1-7 | 93% / 100% | 60% / 69% | 1% / 100% | 79% / 98% | 43% / 100% |
| Random Nums | 6% / 98% | 17% / 29% | 3% / 90% | 1% / 94% | 31% / 100% |
| Letters A-G | 32% / 92% | 63% / 75% | 15% / 97% | 99% / 100% | 59% / 100% |
| Random Letters | 0% / 83% | 49% / 89% | 7% / 74% | 7% / 82% | 7% / 93% |
| No Annotations | 0% / 99%† | 1% / 1%† | 4% / 14%† | 1%/ 3%† | 22% / 41%† |
| Arrow on Plus: Direct / VC-CoT | |||||
| Numbers 1-15 | 29% / 82% | 0% / 61% | 13% / 92% | 97% / 99% | 88% / 99% |
| Random Nums | 0% / 5% | 0% / 5% | 0% / 0% | 1% / 78% | 3% / 77% |
| Letters A-O | 0% / 3% | 0% / 1% | 0% / 2% | 75% / 74% | 14% / 99% |
| Random Letters | 0% / 0% | 0% / 2% | 0% / 1% | 0% / 25% | 3% / 96% |
| No Annotations | 0% / 0%† | 0% / 0%† | 0% / 0%† | 0% / 1%† | 0% / 1%† |
@@
Table 4 shows the performance of Qwen2-VL, Molmo, Janus-Pro, GPT-4-Turbo, and GPT-4o, and Appendix Section G shows results for all other models. Molmo and Janus-Pro were explicitly designed for CoT reasoning through dedicated training, while GPT models are well-documented to excel with CoT prompting Wei et al. (2023), making them the most relevant candidates for VC-CoT. In Table 4, we see that these models exhibit substantial improvements when visually cued CoT prompts are paired with explicit visual annotations, such as numbering or labeling each side. This is especially evident in the GPT models, where GPT-4o increases from 7% to 93%, in the random letters annotation.
@@ Appendix H VC-CoT Results on MathVerse
Despite these challenges, our results show that VC-CoT improves overall performance across these models, with an average accuracy increase from 34.51% (Direct) to 37.86% (VC-CoT). Notably, 9 out of 13 models (including ones in Table 5) benefit from visually guided reasoning, suggesting that explicit extraction and processing of visual information can help compensate for the weaknesses of generic CoT prompting.
@@
| Numbers 1-7 | 0% | 1% | 0% | 5% | 0% | 17% | 93% | 100% | 26% | 68% | 100% | 100% | 0% | 0% | 79% | 81% | 0% | 0% |
| Random Numbers | 0% | 0% | 0% | 0% | 0% | 0% | 6% | 98% | 1% | 21% | 32% | 50% | 0% | 0% | 2% | 0% | 0% | 1% |
| Letters A-G | 0% | 0% | 0% | 0% | 0% | 0% | 32% | 92% | 1% | 2% | 45% | 96% | 0% | 0% | 0% | 0% | 0% | 0% |
| Random Letters | 0% | 0% | 0% | 0% | 0% | 0% | 0% | 83% | 0% | 3% | 3% | 58% | 0% | 0% | 0% | 0% | 0% | 1% |
| No Annotations | 0% | 0% | 0% | 0% | 0% | 0% | 12% | 99% | 0% | 0% | 0% | 0% | 0% | 0% | 0% | 0% | 0% | 0% |
| Numbers 1-15 | 0% | 8% | 12% | 28% | 26% | 26% | 29% | 82% | 0% | 14% | 99% | 99% | 0% | 0% | 0% | 0% | 0% | 61% |
| Random Numbers | 0% | 0% | 0% | 0% | 0% | 0% | 0% | 5% | 0% | 3% | 32% | 21% | 0% | 0% | 0% | 0% | 0% | 28% |
| Random Letters | 0% | 0% | 0% | 0% | 0% | 0% | 0% | 0% | 0% | 1% | 0% | 84% | 0% | 0% | 0% | 0% | 0% | 0% |